OpenStudy (anonymous):

A woman drives 10 miles, accelerating uniformly from rest to 70 mph. How long does it take for her to reach 70 mph? in hours

OpenStudy (anonymous):

Recall v = (d/t) Where: v = velocity d = distance t = time Therefore: 70 = 10/t t = 10/70